OpenStudy (anonymous):

compute J=∫∫√(x^2+y^2)dx dy. x^2+y^2≤4. i've convert the integral to polar coordinates and i've got ∫∫ρ2dθdρ. but i'm stuck. how do i convert x^2+y^2≤4?

OpenStudy (loser66):

 any idea?

OpenStudy (anonymous):

nope

OpenStudy (loser66):

why? you study double integral, at least, you have "something" to solve the problem, right? Cannot solve if you don't know anything, If I work by myself, how can you understand mine with that "nope"?

zepdrix (zepdrix):

Hmm have you tried converting to polar? :)

OpenStudy (anonymous):

no i didnt. i'll try now

zepdrix (zepdrix):

I think it will work out really nicely. Lemme know if you get stuck.

OpenStudy (anonymous):

i dont have the answer for the problem to compare. for the second integral i noticed from the book that it should have 0 and pi buy i don't understand how the first one needs to be. can you please explain it to me? i've found more examples and neither had a pattern that i could follow

OpenStudy (john_es):

As you have said, you have the condition, \[x^2+y^2\leq 4\] When you apply the polar change, then, it should be, \[0\leq \rho^2\leq 4\Rightarrow 0\leq \rho\leq 2\] So (now corrected) \[\int_0^2\int_0^{2\pi}\rho^2d\rho d\theta=\frac{2\pi}{3}\rho^3=\frac{16\pi}{3}\]

OpenStudy (anonymous):

thank you so much.